40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 293     Entry time: Fri Feb 1 17:17:05 2008
Author: John 
Type: Summary 
Category: PSL 
Subject: Mach-Zender tweaking 
I helped Rob adjust the alignment of the Mach-Zender to try and reduce any AM on the laser light. Our goal was to reduce the large offsets in the DD signals.

We reduced the MZ refl from 0.54 to 0.39. We were able to re-lock the mode cleaner without problems. We then centred the WFS heads.

No change in the DD signal offsets.
ELOG V3.1.3-